In a titration, 25.00mL of 0.10mol L-1 sodium hydroxide solution was required to neutralize 12.5 mL of hydrochloric acid. Calculate the concentration of hydrochloric acid.
1
Expert's answer
2019-05-01T04:28:10-0400

0.0025mol   X mol

    NaOH    +   HCl = NaCl + H2O

    1 mol       1 mol


n(NaOH) = C(NaOH) × V(NaOH) = 0.10 mol L-1 × 0.025L = 0.0025 mol.

n(HCl) = X mol = 0.0025 mol.

C(HCl) = n(HCl) / V(HCl) = 0.0025 mol / 0.0125 L = 0.20 mol L-1.